OpenStudy (anonymous):

(A) Show 10/20 and 30/60 form a proportion by finding a common multiplier. (B)Show that ratios in part (a) are equal by writing them in simplest form.

OpenStudy (amonoconnor):

Those aren't equal though... 10/30 = 1/3 = .33333... (and on an on) 30/60= 3/6 = 1/2 = .5 Sorry!

OpenStudy (amonoconnor):

Everything can be divided by the Least Common Factor of 2, but that just makes everything smaller... Not equal :/

OpenStudy (amonoconnor):

There is no "common multiplier" that when multiplied with both 10 and 30, creates 30 and 60. Are you sure you typed the right numbers in your question?

OpenStudy (anonymous):

@amonoconnor

OpenStudy (amonoconnor):

Yes? :)

OpenStudy (anonymous):

Can you help me do this question? @amonoconnor

OpenStudy (amonoconnor):

Umm, yes. The one above you mean?

OpenStudy (anonymous):

Can you help me with this question?